A company with working capital of $397,337 and a current ratio of 2.8 pays a $76,424 short-term liability. Determine the amount of working capital immediately after payment.

Select the correct answer.

$473,761

$320,913

$76,424

$397,337

2)

Based on the following data for the current year, what is the number of days' sales in receivable?

Net sales on account during year $528,447
Cost of goods sold during year 190,300
Accounts receivable, beginning of year 48,837
Accounts receivable, end of year 52,362
Inventory, beginning of year 87,713
Inventory, end of year 118,599

Round your answer up to the nearest whole day.

Select the correct answer.

82 days

71 days

35 days

131 days

3)

Based on the following data for Privett Company, what is the amount of working capital?

Privett Company
Accounts payable $31,171
Accounts receivable 60,469
Accrued liabilities 6,767
Cash 20,739
Intangible assets 37,411
Inventory 80,015
Long-term investments 103,059
Long-term liabilities 79,111
Marketable securities 35,033
Notes payable (short-term) 29,984
Property, plant, and equipment 608,756
Prepaid expenses 1,697

Select the correct answer.

$947,179

$116,241

$130,031

$197,953
Answer #1

1) Working capital = Current assets-Current liabilities

397337 = 2.8X-X

X = 397337/1.8 = 220743

Current assets = 220743*2.8 = 618080

Working Capital = (618080-76424)-(220743-76424) = 397337

So answer is d) $397337

2) Average account receivable = (48837+52362/2) = 50599.50

Number of sales in receivable = 50599.5*365/528447 = 34.95 Days

So answer is c) 35 Days

3) Current assets = 60469+20739+80015+35033+1697 = 197953

Current liabilities = 31171+6767+29984 = 67922

Working Capital = 197953-67922 = 130031

So answer is c) $130031